# Positioning-Aided Channel Estimation for Multi-LEO Satellite Downlink Communications

### v1

@@ III Two-timescale Positioning-Aided Channel Estimation
To characterize the performance limits of the proposed two-timescale PACE, we conduct a two-phase theoretical performance analysis. In the first phase, we analyze the downlink multi-LEO satellite-based positioning performance, while in the second phase, we assess the uplink channel estimation performance.
@@
Before analyzing positioning performance, we acknowledge the presence of model mismatches due to hardware imperfections. First, low-cost oscillators and inherent hardware defects can introduce phase and frequency errors [32]. Second, clock biases are inevitable between satellites and UTs as well as between satellites themselves. Although a sub-nanosecond clock synchronization can be achieved by deploying GNSS receivers on LEO satellites [33], actual satellite signals remain subject to ionospheric and tropospheric delays during propagation [34]. These delays, influenced by factors such as temperature, pressure, humidity, and the satellite’s elevation angle relative to the UT, result in unique delays for each satellite-UT pair, effectively introducing an additional clock bias in the signal model. Despite these non-idealities, many localization algorithms simplify the model by ignoring such factors to reduce computational complexity. For instance, in the scenario considered here, the large beam footprint of LEO satellite communication systems relaxes UT positioning accuracy requirements, while rapid channel variations prioritize model simplicity for timely estimations. Accordingly, we adopt a simplified model to estimate unknowns derived from a more complex realistic model, which inherently leads to model mismatches. Similarly, in the second phase, namely position information-aided channel estimation, the presence of estimation errors in UT positions also introduces model mismatches. Consequently, the performance bound under these model mismatches can be effectively analyzed using the tool of MCRB [27].
III-A MCRB Basics
This subsection briefly recaps the fundamentals of MCRB. For a parameter $\bar{\bm{\theta}}\in\mathbb{R}^{M\times 1}$, the lower bound (LB) matrix of the mean squared error of a mismatched estimator is provided by[27]
where $\mathcal{D}\{\xi_{\text{T}}\left(\mathbf{y};\bm{\theta}\right)\left|\right|\xi_{\text{M}}(\mathbf{y};\tilde{\bm{\theta}})\}$ denotes the Kullback-Leibler (KL) divergence between the true likelihood function $\xi_{\text{T}}\left(\mathbf{y};\bm{\theta}\right)$ and the mismatched likelihood function $\xi_{\text{M}}(\mathbf{y};\tilde{\bm{\theta}})$. In addition, $\mathbf{A}_{\tilde{\bm{\theta}}}$ and $\mathbf{B}_{\tilde{\bm{\theta}}}$ represent two generalized Fisher information matrixs, whose elements in the $i$-th row and the $j$-th column are determined by[27]
and
respectively. Here, $\mathbf{C}_{\text{M}}$ represents the covariance matrix of $\xi_{\text{M}}(\mathbf{y};\bm{\theta})$ and $\bm{\epsilon}(\tilde{\bm{\theta}})=\bm{\kappa}(\bar{\bm{\theta}})-\bm{\mu}(\tilde{\bm{\theta}})$ with $\bm{\kappa}(\bar{\bm{\theta}})$ and $\bm{\mu}(\tilde{\bm{\theta}})$ being the noise-free observations under the true and mismatched models, respectively[27, 35].
III-B Large-Timescale Downlink Positioning
III-B1 Signal Model
@@
Let $\mathbf{t}_{s}[\ell,k]\in\mathbb{C}^{U}$ be the pilot (with unit-modulus elements) sent by the $s$-th LEO satellite over the $\ell$-th symbol and the $k$-th subcarrier. We assume the satellites are multiplexed in the positioning phase through time-division multiple access (TDMA)[25]. The received baseband signal at the $u$-th user, from the $s$-th satellite can be expressed as
@@
$=$ $\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\mathbf{a}^{\mathsf{T}}\left(\bm{\theta}_{s,u}\right)\mathbf{f}_{s}\left[\ell,k\right]+z_{u}\left[\ell,k\right],$ (9)
@@
Let ${\bm{\eta}}_{s,u}=[\acute{\bm{\eta}}_{s,u}^{\mathsf{T}},\grave{\bm{\eta}}_{s,u}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{6}$ be the unknown channel-domain parameters in signals received from the $s$-th satellite, where $\acute{\bm{\eta}}_{s,u}=[\upsilon_{s,u},\tau_{s,u}]^{\mathsf{T}}\in\mathbb{R}^{2}$ contains the channel-domain parameters we used to localize the user and $\grave{\bm{\eta}}_{s,u}=[\bm{\theta}_{s,u}^{\mathsf{T}},\Re(\alpha_{s,u}),\Im(\alpha_{s,u})]^{\mathsf{T}}\in\mathbb{R}^{4}$ contains the other nuisance parameters that we do not explore. Note here the AOD $\bm{\theta}_{s,u}$ is unused for UT positioning due to the limited angular resolution of satellites. We further concatenate all the utilized channel-domain parameters collected from $S$ satellites as $\acute{\bm{\eta}}_{u}=[\acute{\bm{\eta}}_{1,u}^{\mathsf{T}},\ldots,\acute{\bm{\eta}}_{S,u}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{2S}$. Therefore, in the large-timescale downlink positioning, we first estimate $\acute{\bm{\eta}}_{u}$ based on $y_{s,u}[\ell,k]$, $s=1,\dots,S$, $\ell=1,\dots,L_{\text{p}}$, $k=1,\dots,K$, where $L_{\text{p}}$ denotes the number of pilot symbols for downlink positioning. Afterwards, we estimate the $u$-th user’s position, clock biases, and CFOs.
III-B2 The FIM of ${\bm{\eta}}_{s,u}$
@@
Let ${\mu}_{s,u}[\ell,k]$ denote the noise-free version of $y_{s,u}[\ell,k]$ and $\mathbf{G}_{s,u}[\ell,k]=[\mathbf{g}_{v}[\ell,k],\mathbf{g}_{\tau}[\ell,k],\mathbf{g}_{\theta_{\mathrm{az}}}[\ell,k],\mathbf{g}_{\theta_{\mathrm{el}}}[\ell,k],\mathbf{g}_{\alpha,\mathrm{R}}[\ell,k],\mathbf{g}_{\alpha,\mathrm{I}}[\ell,k]]\in\mathbb{C}^{N\times 6}$. Here, the columns fulfill
$\mathbf{g}_{v}\left[\ell,k\right]$ $=\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\left(\jmath 2\pi\ell T\right)\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (10a)
@@
$\mathbf{g}_{\tau}\left[\ell,k\right]$ $=\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\left(-\jmath 2\pi k\Delta f\right)\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (10b)
$\mathbf{g}_{\theta_{\mathrm{az}}}\left[\ell,k\right]$ $=\alpha_{s,u}\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\frac{\partial\mathbf{a}\left(\bm{\theta}_{s,u}\right)}{\partial\theta_{\mathrm{az}}},$ (10c)
$\mathbf{g}_{\theta_{\mathrm{el}}}\left[\ell,k\right]$ $=\alpha_{s,u}\left(\frac{\partial\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]}{\partial\theta_{s,u}^{\mathrm{el}}}\mathbf{a}\left(\bm{\theta}_{s,u}\right)\right.$ (10d)
@@
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left.\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\frac{\partial\mathbf{a}\left(\bm{\theta}_{s,u}\right)}{\partial\theta_{s,u}^{\mathrm{el}}}\right),$
@@
$\mathbf{g}_{\alpha,\mathrm{R}}\left[\ell,k\right]$ $=\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (10e)
@@
$\mathbf{g}_{\alpha,\mathrm{I}}\left[\ell,k\right]$ $=\jmath\gamma_{\theta_{s,u}^{\mathrm{el}}}\left[\ell,k\right]\mathbf{a}\left(\bm{\theta}_{s,u}\right).$ (10f)
@@
Based on the signal received from the $s$-th satellite, the FIM of ${\bm{\eta}}_{s,u}$ can be computed by the Slepian-Bangs formula as
@@ III-B3 The FIM of $\acute{\bm{\eta}}_{u}$
Based on this, we can infer that the achievable lowest estimation error on $\acute{\bm{\eta}}_{u}$ regardless of the adopted algorithms is characterized by the Gaussian distribution $\mathcal{N}\big{(}\mathbf{0},\mathbf{J}_{s}^{-1}(\acute{\bm{\eta}}_{u})\big{)}$. Any algorithm achieving this error covariance matrix is called an efficient estimator.
III-B4 MCRB and LB Derivation
@@
Let $\mathbf{v}_{s}\in\mathbb{R}^{3}$ be the velocity of the $s$-th LEO satellite. Here, as the velocity of LEO satellite is usually much higher than that of the UT, we ignore the velocity contribution from the UT side. Based on the underlying geometric relationship, the forward model between the entries of $\acute{\bm{\eta}}_{u}$ and unknown position of the $u$-th user can be expressed as
$\upsilon_{s,u}$ $=\frac{\mathbf{v}_{s}^{\mathrm{T}}(\mathbf{p}_{u}-\mathbf{p}_{s})}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}+\delta_{u}+\Delta\delta_{s,u},$ (13a)
$\tau_{s,u}$ $=\frac{\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}{c}+b_{u}+\Delta b_{s,u},$ (13b)
which we refer to as the true model. In real applications, we often ignore these clock bias and CFO differences, i.e., assuming all satellite-UT pairs have the same clock bias and CFO, for algorithmic simplicity. To account for the impact of such model mismatch, here we derive the MCRB. When neglecting the difference on each clock bias and CFO, the mismatched model is given by
$\upsilon_{s,u}$ $=\frac{\mathbf{v}_{s}^{\mathrm{T}}(\mathbf{p}_{u}-\mathbf{p}_{s})}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}+\delta_{u},$ (14a)
$\tau_{s,u}$ $=\frac{\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}{c}+b_{u}.$ (14b)
Now, we suppose a mismatched estimator that aims to estimate the unknown position of the $u$-th user based on the realistic observation $\acute{\bm{\eta}}_{u}$ and simplified model (14) . the total unknown parameters in the mismatch model can be concatenated as $\mathbf{r}_{u}=[\mathbf{p}_{u}^{\mathsf{T}},\delta_{u},b_{u}]^{\mathsf{T}}\in\mathbb{R}^{5}$. Based on the true model (13) and mismatched model (14), we can respectively express the true likelihood function $\xi_{\mathrm{T}}$ and the mismatched likelihood function $\xi_{\mathrm{M}}$ as
$\ln{\xi_{\mathrm{T}}}\left(\acute{\bm{\eta}}_{u};\mathbf{r}_{u}\right)\propto\big{(}\acute{\bm{\eta}}_{u}-\bar{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)}^{\mathsf{T}}\bm{\Sigma}^{-1}\big{(}\acute{\bm{\eta}}_{u}-\bar{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)},$ (15)
@@
$\ln{\xi_{\mathrm{M}}}\left(\acute{\bm{\eta}}_{u};\mathbf{r}_{u}\right)\propto\big{(}\acute{\bm{\eta}}_{u}-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)}^{\mathsf{T}}\bm{\Sigma}^{-1}\big{(}\acute{\bm{\eta}}_{u}-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big{)},$ (16)
respectively, where the function $\bar{\mathbf{f}}(\cdot)$ maps $\mathbf{r}_{u}$ to $\acute{\bm{\eta}}_{u}$ according to (13), and the function $\tilde{\mathbf{f}}(\cdot)$ maps $\mathbf{r}_{u}$ to $\acute{\bm{\eta}}_{u}$ according to (14). As mentioned in Section III-B3, $\acute{\bm{\eta}}_{u}$ is the estimated channel-domain parameters through an efficient estimator, i.e., $\acute{\bm{\eta}}_{u}\sim\mathcal{N}(\bar{\mathbf{f}}(\mathbf{r}_{u}),\bm{\Sigma})$, where $\bm{\Sigma}=\mathbf{J}^{-1}(\acute{\bm{\eta}}_{u})$.
@@
  • MCRB: According to (III-A) and (III-A), the generalized FIMs, i.e., $\mathbf{A}_{\tilde{\mathbf{r}}_{u}}$ and $\mathbf{B}_{\tilde{\mathbf{r}}_{u}}$, can be computed by determining the first- and second-order derivatives of $\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)$ with respect to the elements of $\mathbf{r}_{u}$, as detailed in Appendix A. Subsequently, the MCRB is expressed as
@@
  • LB: Finally, the estimation error LB matrix can be computed according to (5) as
@@
III-C Small-Timescale Uplink Channel Estimation
III-C1 Signal Model

### v4

@@ III Two-timescale Positioning-Aided Channel Estimation
To characterize its performance limits, we conduct a two-phase theoretical performance analysis [^5] For both downlink positioning and uplink channel estimation, we theoretically characterize the performance under model mismatch using the MCRB, with the aim of revealing the potential of the proposed PACE framework in enhancing channel estimation. The design of specific estimators is beyond the scope of this paper and is left for future work. for PACE. In the first phase, we analyze the downlink multi-LEO satellite-based positioning performance, while in the second phase, we assess the uplink channel estimation performance.
@@
Note that model mismatches can arise from various hardware imperfections. [^6] While impairments such as phase noise and power amplifier nonlinearities are also relevant at Ku/Ka band[47], we focus on clock bias and CFO because they directly perturb delay and Doppler, introducing deterministic biases across satellites that dominate positioning errors. By contrast, phase noise and nonlinearities mainly act as stochastic distortions that reduce effective received power and can, in principle, be mitigated on a single-satellite basis. First, clock biases are inevitable between satellites and UTs as well as among satellites themselves. Although sub-nanosecond synchronization can be achieved by equipping LEO satellites with GNSS receivers [48], the received signals remain subject to ionospheric and tropospheric delays during propagation [49]. These delays, shaped by temperature, pressure, humidity, and the satellite’s elevation angle relative to the UT, induce unique offsets for each satellite–UT pair and effectively translate into equivalent clock biases in the signal model. Second, low-cost oscillators and inherent hardware impairments can introduce CFO [50]. Many localization algorithms simplify the underlying model by neglecting certain factors to reduce analytical complexity. To capture the impact of this simplification, in the positioning phase we employ a simplified model to estimate unknowns that are actually generated by a more realistic one, which inherently introduces model mismatches. Similarly, in the second phase, position information-aided channel estimation, estimation errors in UT positions also lead to mismatches. Consequently, the performance bounds under such conditions can be effectively characterized using the MCRB [38, 51].
III-A Large-Timescale Downlink Positioning
III-A1 Signal Model
@@
Let $\mathbf{t}_{s}[\ell,k]\in\mathbb{C}^{U}$ be the pilot (with unit-modulus elements) sent by the $s$-th LEO satellite over the $\ell$-th symbol and the $k$-th subcarrier. We assume the satellites are multiplexed in the positioning phase through time-division multiple access (TDMA)[36]. The received baseband signal at the $u$-th UT, from the $s$-th satellite can be expressed as
@@
$=$ $\vartheta_{s,u}\gamma\left[\ell,k\right]\mathbf{a}^{\mathsf{T}}\left(\bm{\theta}_{s,u}\right)\mathbf{f}_{s}\left[\ell,k\right]+z_{u}\left[\ell,k\right],$ (5)
@@
Let ${\bm{\eta}}_{s,u}=[\acute{\bm{\eta}}_{s,u}^{\mathsf{T}},\grave{\bm{\eta}}_{s,u}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{6}$ denote the unknown channel-domain parameters in the signal received from the $s$-th satellite, where $\acute{\bm{\eta}}_{s,u}=[\upsilon_{s,u},\tau_{s,u}]^{\mathsf{T}}\in\mathbb{R}^{2}$ are the parameters used for UT positioning, and $\grave{\bm{\eta}}_{s,u}=[\bm{\theta}_{s,u}^{\mathsf{T}},\Re(\vartheta_{s,u}),\Im(\vartheta_{s,u})]^{\mathsf{T}}\in\mathbb{R}^{4}$ are treated as nuisance parameters. Delay $\tau_{s,u}$ and Doppler $\upsilon_{s,u}$ are directly determined by the UT’s position and velocity relative to the satellites, making them the most reliable observables for positioning. In contrast, the AOD $\bm{\theta}_{s,u}$ is not exploited here due to the limited angular resolution of LEO satellites operating at high altitudes, which makes AOD estimates unreliable unless extremely large arrays are employed. Similarly, the channel gain $\vartheta_{s,u}$ is heavily influenced by environmental variability and hardware impairments and thus provides little stable geometric information for positioning, except in fingerprinting-based approaches [40].
@@
We further concatenate all the utilized channel-domain parameters collected from $S$ satellites as $\acute{\bm{\eta}}_{u}=[\acute{\bm{\eta}}_{1,u}^{\mathsf{T}},\ldots,\acute{\bm{\eta}}_{S,u}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{2S}$. Therefore, in the large-timescale downlink positioning, we first estimate $\acute{\bm{\eta}}_{u}$ based on $y_{s,u}[\ell,k]$, $s=1,\dots,S$, $\ell=1,\dots,L_{\text{p}}$, $k=1,\dots,K$, where $L_{\text{p}}$ denotes the number of pilot symbols for downlink positioning. Afterwards, we estimate the $u$-th UT’s position, clock biases, and CFOs.
III-A2 The FIM of ${\bm{\eta}}_{s,u}$
@@
Let ${\mu}_{s,u}[\ell,k]$ denote the noise-free version of $y_{s,u}[\ell,k]$ and $\mathbf{G}_{s,u}[\ell,k]=[\mathbf{g}_{v}[\ell,k],\mathbf{g}_{\tau}[\ell,k],\mathbf{g}_{\theta_{\mathrm{az}}}[\ell,k],\mathbf{g}_{\theta_{\mathrm{el}}}[\ell,k],\mathbf{g}_{\vartheta,\mathrm{R}}[\ell,k],\mathbf{g}_{\vartheta,\mathrm{I}}[\ell,k]]\in\mathbb{C}^{N\times 6}$. Here, the columns fulfill
$\mathbf{g}_{v}\left[\ell,k\right]$ $=\vartheta_{s,u}\gamma\left[\ell,k\right]\left(\jmath 2\pi\ell T\right)\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (6a)
@@
$\mathbf{g}_{\tau}\left[\ell,k\right]$ $=\vartheta_{s,u}\gamma\left[\ell,k\right]\left(-\jmath 2\pi k\Delta f\right)\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (6b)
@@
$\mathbf{g}_{\theta_{\mathrm{az}}}\left[\ell,k\right]$ $=\vartheta_{s,u}\gamma\left[\ell,k\right]\frac{\partial\mathbf{a}\left(\bm{\theta}_{s,u}\right)}{\partial\theta_{s,u}^{\mathrm{az}}},$ (6c)
@@
$\mathbf{g}_{\theta_{\mathrm{el}}}\left[\ell,k\right]$ $=\vartheta_{s,u}\gamma\left[\ell,k\right]\frac{\partial\mathbf{a}\left(\bm{\theta}_{s,u}\right)}{\partial\theta_{s,u}^{\mathrm{el}}},$ (6d)
$\mathbf{g}_{\vartheta,\mathrm{R}}\left[\ell,k\right]$ $=\gamma\left[\ell,k\right]\mathbf{a}\left(\bm{\theta}_{s,u}\right),$ (6e)
@@
$\mathbf{g}_{\vartheta,\mathrm{I}}\left[\ell,k\right]$ $=\jmath\gamma\left[\ell,k\right]\mathbf{a}\left(\bm{\theta}_{s,u}\right).$ (6f)
@@
Based on the signal received from the $s$-th satellite, the Fisher information matrix (FIM) of ${\bm{\eta}}_{s,u}$ can be computed by the Slepian-Bangs formula as
@@ III-A3 The FIM of $\acute{\bm{\eta}}_{u}$
Based on this, we can infer that the achievable lowest estimation error on $\acute{\bm{\eta}}_{u}$ regardless of the adopted algorithms is characterized by the Gaussian distribution $\mathcal{N}\big(\mathbf{0},\mathbf{J}^{-1}(\acute{\bm{\eta}}_{u})\big)$. Any algorithm achieving this error covariance matrix is called an efficient estimator.
III-A4 MCRB and LB Derivation
@@
Let $\mathbf{v}_{s}\in\mathbb{R}^{3}$ be the velocity of the $s$-th LEO satellite. Here, as the velocity of LEO satellite is usually much higher than that of the UT, we ignore the velocity contribution from the UT side. This assumption is justified since LEO satellites move at around 7.6 km/s, whereas UTs typically move at only a few m/s (pedestrians) to a few hundred m/s (vehicles/aircraft), making their velocity contribution two to three orders of magnitude smaller and thus negligible for the Doppler term. Based on the underlying geometric relationship, the forward model between the entries of $\acute{\bm{\eta}}_{u}$ and unknown position of the $u$-th UT can be expressed as
$\upsilon_{s,u}=\frac{\mathbf{v}_{s}^{\mathrm{T}}(\mathbf{p}_{u}-\mathbf{p}_{s})}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}+\delta_{u}+\Delta\delta_{s,u}$ (9)
@@
$\tau_{s,u}=\frac{\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}{c}+b_{u}+\Delta b_{s,u},$ (10)
which we refer to as the true model. Here, $\mathbf{p}_{u}\in\mathbb{R}^{3}$ and $\mathbf{p}_{s}\in\mathbb{R}^{3}$ represent the positions of the $u$-th UT and the $s$-th satellite [^7] Note that LEO satellites’ positions are assumed precisely known at the UTs. This information can be obtained from the ephemerides contained in the two-line element files, which are updated daily by the North American Aerospace Defense Command (NORAD) [36]., respectively.
In real applications, we often ignore these clock bias and CFO differences, i.e., assuming all satellite-UT pairs have the same clock bias and CFO, for algorithmic simplicity. To account for the impact of such model mismatch, we derive the MCRB, with its fundamentals detailed in Appendix A for completeness. When neglecting the difference on each clock bias and CFO, the mismatched model is given by
$\upsilon_{s,u}=\frac{\mathbf{v}_{s}^{\mathrm{T}}(\mathbf{p}_{u}-\mathbf{p}_{s})}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}+\delta_{u}$ (11)
and
$\tau_{s,u}=\frac{\|\mathbf{p}_{u}-\mathbf{p}_{s}\|}{c}+b_{u}.$ (12)
Now, we suppose an estimator that aims to estimate the unknown position of the $u$-th UT based on the realistic observation $\acute{\bm{\eta}}_{u}$ and mismatched model, the total unknown parameters in the mismatch model can be concatenated as $\mathbf{r}_{u}=[\mathbf{p}_{u}^{\mathsf{T}},\delta_{u},b_{u}]^{\mathsf{T}}\in\mathbb{R}^{5}$. Based on the true model and mismatched model, we can respectively express the true likelihood function $\xi_{\mathrm{T}}$ and the mismatched likelihood function $\xi_{\mathrm{M}}$ as
${\xi_{\mathrm{T}}}\left(\acute{\bm{\eta}}_{u};\mathbf{r}_{u}\right)\propto e^{\big(\acute{\bm{\eta}}_{u}-\bar{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big)^{\mathsf{T}}\bm{\Sigma}^{-1}\big(\acute{\bm{\eta}}_{u}-\bar{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big)}$ (13)
and
${\xi_{\mathrm{M}}}\left(\acute{\bm{\eta}}_{u};\mathbf{r}_{u}\right)\propto e^{\big(\acute{\bm{\eta}}_{u}-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big)^{\mathsf{T}}\bm{\Sigma}^{-1}\big(\acute{\bm{\eta}}_{u}-\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)\big)},$ (14)
[…]
@@
  • MCRB: According to Appendix A, the generalized FIMs, i.e., $\mathbf{A}_{\tilde{\mathbf{r}}_{u}}$ and $\mathbf{B}_{\tilde{\mathbf{r}}_{u}}$, can be computed by determining the first- and second-order derivatives of $\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)$ with respect to the elements of $\mathbf{r}_{u}$, as detailed in Appendix B. Subsequently, the MCRB is expressed as
@@
  • LB: Finally, the estimation error lower bound (LB) matrix can be computed as
@@
III-B Small-Timescale Uplink Channel Estimation
III-B1 Signal Model